Solve for x. Give exact answer.
log(x + 2) - log x = 1
X=?

Answer by nerdybill(6948)

log(x + 2) - log x = 1
log(x + 2)/x = 1
(x + 2)/x = 10^1
(x + 2)/x = 10
x + 2 = 10x
2 = 9x
2/9 = x